OpenStudy (anonymous):

A box contains 2 strawberry yogurts, 4 vanilla yogurts and 6 cherry yogurts. Three yogurts are selected at random from the box. Calculate the probability that at least one of the selected yogurts is a cherry yogurt.

OpenStudy (anonymous):

fraction decimal or percent?

OpenStudy (anonymous):

1/2 .5 50%

OpenStudy (anonymous):

6/12 =1/2. There's a 50% probability that a cherry yogurt will be selected. Fraction: 1/2 Decimal: .50 or .5 Percent: 50%

OpenStudy (anonymous):

is the probability 12 yogerts 6 are cherry 6/12 1/2 .5 50%

OpenStudy (anonymous):

glad i could help!

OpenStudy (zarkon):

is the selection of the 3 with or without replacement

OpenStudy (zarkon):

you figure out if it is with/without replacement...then calculate (probability that at least one of the selected yogurts is a cherry yogurt.) =1-(probability that none of the selected yogurts is a cherry yogurt.)

OpenStudy (anonymous):

Without replacement.

OpenStudy (zarkon):

then the prob you select none of the cherry is \[\frac{6}{12}\frac{5}{11}\frac{4}{10}=\frac{1}{11}\] thus your probability is \[1=\frac{1}{11}=\frac{10}{11}\]

OpenStudy (zarkon):

typo \[1-\frac{1}{11}=\frac{10}{11}\]

OpenStudy (anonymous):

That's not probability

OpenStudy (zarkon):

?

OpenStudy (anonymous):

Thanks anyway guys :)!

OpenStudy (zarkon):

the answer is 10/11

OpenStudy (zarkon):

this type of problem follows the hypergeometric distribution. you can write the above as \[1-\frac{{6\choose 3}}{{12\choose3}}=\frac{10}{11}\]

OpenStudy (zarkon):

you can also write it directly .. \[\sum_{k=1}^{3}\frac{{6\choose k}{6\choose 3-k}}{{12 \choose 3}}=\frac{10}{11}\]

OpenStudy (anonymous):

Thanks Zarkon, but thats the wrong way, Thanks anyway :)

OpenStudy (zarkon):

it is not.